The product of 3 more than a number and 11 is 165. What is the number?
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

(x+3)(11)=165
11x+33+165
11x=165-33
11x=132
x=132/11
x=12

Answer by stanbon(75887)

The product of 3 more than a number and 11 is 165. What is the number?
------
Equation:
(x+3)*11 = 165
---
x+3 = 165/11 = 15
----
x = 12
==========================
cheers,
Stan H.
=============